OpenStudy (anonymous):

if h(x)=(fog)(x) and h(x)= 3 root x+3, find g(x) if f(x) = 3 root x+2 thanks

OpenStudy (anonymous):

the 3 would go alittle high to the left of the square root sign i hope you understand

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@danish071996

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

fog means f times g sorry

OpenStudy (dumbcow):

\[\sqrt[3]{g(x) +2} = \sqrt[3]{x+3}\] raising both sides to 3rd power cancels out the 3rd roots \[g(x) +2 = x+3\] \[g(x) = x+1\]